OpenStudy (miss.rose):

I'll Fan and Medal!!! PLEASE HELP!!! A line goes through the point (–2,3) and has a slope of –4. What is the y-intercept of the line?

OpenStudy (mrnood):

So you must be studying the slope/intercept equation of a line you recognise y = mx + b ?

OpenStudy (miss.rose):

yes. I just don't know how to set it up.

OpenStudy (mrnood):

so - what does the value of m represent in that equation?

OpenStudy (miss.rose):

not. sure. I missed like 3 days of the class, because I was sick. And the notes my friend gave me make no sense.

OpenStudy (mrnood):

this is a basic equation that you will need for nearly all questions on straight lines. It is important that you learn it for yourself - if I just tell oyu it won't sink in. Look here and find out what m is in the equation y=mx+b http://www.coolmath.com/algebra/08-lines/07-equations-of-lines-slope-intercept-form-01

OpenStudy (miss.rose):

m represents slope.

OpenStudy (mrnood):

ok good now the question tells you that the slope is -4 so we can begin to write the equation y= -4x+b OK so far?

OpenStudy (miss.rose):

Yes.

OpenStudy (mrnood):

so you know the slope and you know that the point (–2,3) lies on the line so when x = -2 y=3 so can oyu write my equation from above, but substitute 3 for y , and -2 for x

OpenStudy (miss.rose):

y=mx+b 3=(-4)(-2)+b What's b?

OpenStudy (mrnood):

b is what you are trying to find (the intercept) so if we tidy up the equation we get 3= 8 +b so subtract 8 from each side to find b

OpenStudy (miss.rose):

b=-5

OpenStudy (mrnood):

good so the total equation is this y= -4x-5 and the graph looks like this:  well done